OpenStudy (anonymous):

Please help, log4(x-2) + 2log4(x) = 1 + log4(2x) Thanks

OpenStudy (saifoo.khan):

is 4 the base?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

log 4 [2x(x)/(2x)] =1 ===> x=4

OpenStudy (anonymous):

log4(x-2) + log4 (x)^2 - log4(2x) = 1 log4[ (x^2(x-2)) / 2x] = log4 4 (x^2(x-2)) / 2x] = 4 x(x-2) = 8 x^2 - 2x - 8 = 0 (x -4)(x+2) = 0 x = 4 or -2

OpenStudy (anonymous):

-2 is not valid x = 4 good work raheed - i did it the long way!!

OpenStudy (anonymous):

awesome, thanks a lot i appreciate it

OpenStudy (anonymous):

yw